OpenStudy (anonymous):

Help me! Find the slope, if it exists x=-7

OpenStudy (anonymous):

does not exist

OpenStudy (anonymous):

the graph of x = -7 is a vertical line, and vertical lines do not have slopes

OpenStudy (anonymous):

so it would be undefined?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

the slope is undefined. One of the previous answerers erred in that he forgot to use 2 points instead of only one. So, let's take the points (1, -7) and (2, -7). m = (2 - 1) / (-7 - -7) m = 1 / 0 Since you cannnot divide by 0, m = undefined. This is always true of vertical lines. The equation of a vertical line will always look like the one you provided, for example: x = 1 x = 2 x = -5 x = -43575986 x = 0